# Bloch-Beilinson conjectures for Hecke characters and Eisenstein cohomology of Picard surfaces

### v2

@@ 3.3. Bloch-Beilinson conjectures for Hecke characters
Consider the motive $M_{\phi}=(h^{1}(A_{\phi})^{\otimes w},e_{\phi})$ attached to $\phi$ as in the previous subsection. Recall that since the integer $\frac{w+1}{2}$ is critical for $M_{\phi}$, Deligne has used the comparison isomorphism between the Betti and de Rham realizations of the weight $-1$ motive $M_{\phi}(\frac{w+1}{2})$ to attach to it a canonical non-zero period $c^{+}(M_{\phi})\in{\mathbb{R}}$. He conjectured that if the value $L(M_{\phi},\frac{w+1}{2})$ does not vanish, then it should be equal to an algebraic multiple of $c^{+}(M_{\phi})$. He offered no statement when $L(\phi,s)$ has a zero at the central point. In order to express what is believed to happen in the latter case, consider the Chow group $CH^{n}(X)$ of a smooth, projective variety $X$ and its subgroup $CH^{n}(X)_{0}$ formed by cycles homologically equivalent to 0, for $n\in\{0,\dots,\dim X\}$. Bloch and Beilinson conjectured the existence of a canonical, non-degenerate height pairing
$h^{n}_{X}:CH^{n}(X)_{0}\otimes{\mathbb{Q}}\times CH^{\dim X+1-n}(X)_{0}\otimes{\mathbb{Q}}\rightarrow{\mathbb{R}}$
@@
In order to do so, we have to extend the conjecture to the Chow groups of our motive $M_{\phi}$. Recall that for $X,n$ as above, there is a natural action of $CH^{\dim X}(X\times X)$ on $CH^{n}(X)$, preserving $CH^{n}(X)_{0}$. Denote by $\Gamma_{*}$ the action of an element $\Gamma\in CH^{\dim X}(X\times X)$. For such a $\Gamma$, the pairing $h^{n}_{X}$ restricts to a pairing, also conjecturally non-degenerate
$h^{n}_{X,\Gamma}:\Gamma_{*}CH^{n}(X)_{0}\otimes{\mathbb{Q}}\times\Gamma_{*}CH^{\dim X+1-n}(X)_{0}\otimes{\mathbb{Q}}\rightarrow{\mathbb{R}}$
@@
    $\mathop{\rm\tiny{ord}}\nolimits_{s=\frac{w+1}{2}}L(\phi,s)=\dim_{{\mathbb{Q}}}(e_{\phi}e_{1,w})_{*}CH^{\frac{w+1}{2}}(A^{w}_{\phi})_{0}\otimes{\mathbb{Q}}$
@@
    $L^{*}(\phi,\frac{w+1}{2})\sim_{\overline{{\mathbb{Q}}}}c^{+}(M_{\phi})\det(h^{\frac{w+1}{2}}_{A_{\phi}^{w},e_{\phi}e_{1,w}})$
@@
Now consider the Abel-Jacobi map
$e_{\phi}e_{1,w}CH^{\frac{w+1}{2}}(A^{w}_{\phi})_{0}\rightarrow\mathop{\rm Ext}\nolimits^{1}_{\mathop{\rm MHS}\nolimits_{\overline{{\mathbb{Q}}}}}(\mathds{1},e_{\phi}e_{1,w}H^{w}(A^{w}_{\phi})(\frac{w+1}{2}))$
Since $A_{\phi}^{w}$ is defined over a number field, this map is conjectured to be injective ([5, 5.6]). Using the description (25) of $H_{\phi}$, it takes the form
(26) $e_{\phi}e_{1,w}CH^{\frac{w+1}{2}}(A^{w}_{\phi})_{0}\rightarrow\mathop{\rm Ext}\nolimits^{1}_{\mathop{\rm MHS}\nolimits_{\overline{{\mathbb{Q}}}}}(\mathds{1},H_{\phi}(\frac{w+1}{2}))$
Let $\epsilon(\phi)$ be the sign of the functional equation of $\phi$. The conjectural injectivity of the map (26), point (1) of Conjecture 3.2, and the fact that $\epsilon(\phi)=-1$ implies that $L(\phi,s)$ vanishes at the central point, lead to the following (considerably weaker and less precise) conjecture.
@@
Let $\phi$ be an algebraic Hecke character of $F$ of odd weight w. If $\epsilon(\phi)=-1$, then there exists a non-trivial extension
@@
A way of making the above property of being of geometric origin precise is the following. There is a triangulated, $\overline{{\mathbb{Q}}}$-linear tensor category $\mathop{\rm DM}\nolimits_{c}(F)_{\overline{{\mathbb{Q}}}}$ of constructible motives over $F$, containing the category of Chow motives as a full subcategory. It is equipped with a realization functor
$R_{H}:\mathop{\rm DM}\nolimits_{c}(F)_{\overline{{\mathbb{Q}}}}\rightarrow D^{b}(\mathop{\rm MHS}\nolimits_{\overline{{\mathbb{Q}}}})$
Then, we say that an object $E$ of $\mathop{\rm MHS}\nolimits_{\overline{{\mathbb{Q}}}}$ is of geometric origin if it belongs to the full Tannakian subcategory of $\mathop{\rm MHS}\nolimits_{\overline{{\mathbb{Q}}}}$ generated by the cohomology objects of objects belonging to the image of $R_{H}$.
@@ 4.1. Boundary of the Baily-Borel compactification
For facts recalled in this paragraph, we refer to [28, Sects. 3.6-3.7]. Any ${\mathbb{Q}}$-conjugacy class of parabolics in $\mathrm{G}$ provides a set of strata of the boundary of the Baily-Borel compactification of $\mathrm{S}_{\Gamma}$. More precisely, given a representative $\mathrm{P}$ of such a conjugacy class, there is a canonical reductive subgroup $\mathrm{M}$ of the Levi component of $\mathrm{P}$, and a canonical conjugacy class of morphisms
@@ 4.2. Boundary of the Borel-Serre compactification
For facts recalled in this paragraph, we refer to [17, pp. 567-570]. The Borel-Serre compactification $\overline{\mathrm{S}}_{\Gamma}$ of $\mathrm{S}_{\Gamma}$ admits a canonical projection $\pi$ to the Baily-Borel one. Let us describe, in our situation, the fiber of $\pi$ over each cusp $p$ of $\mathrm{S}^{*}_{\Gamma}$. To do so, let us write
@@
Denote by $A_{\mathrm{G}}$ the identity component of the real points of the maximal ${\mathbb{Q}}$-split subtorus $\mathrm{T}^{s}$ of $\mathrm{G}^{\mathop{\rm\tiny{der}}\nolimits}$. From equation (10), we see that there is an isomorphism of Lie groups $\mathrm{T}^{\prime}({\mathbb{R}})\simeq{\mathbb{C}}^{\times}$ which by equation (11) restricts to $A_{\mathrm{G}}\simeq{\mathbb{R}}^{\times}_{>0}.$
Define ${}^{0}T$ as the Lie group $A_{\mathrm{G}}\backslash\mathrm{T}^{\prime}({\mathbb{R}})$. It is isomorphic to $S^{1}$. Denote by $K^{T}_{\infty}$ the intersection with ${}^{0}T$ of a maximal compact subgroup of $\mathrm{G}^{\mathop{\rm\tiny{der}}\nolimits}({\mathbb{R}})$. By conjugating the maximal compact subgroup of equation (18) via the change of basis matrix given in (3), one sees that $K^{T}_{\infty}$ coincides with ${}^{0}T$ itself. On the other hand, denote $\Gamma_{\mathrm{U}}:=\Gamma\cap\mathrm{U}^{\prime}({\mathbb{R}})$. The general theory tells us that the space $\pi^{-1}(p)$ has the structure of a fibration, with base a quotient of ${}^{0}T\backslash K^{T}_{\infty}$ (hence a point in our case), and with fiber diffeomorphic to the nilmanifold
$\mathrm{S}_{\mathrm{U}_{\Gamma}}:=\Gamma_{\mathrm{U}}\backslash\mathrm{U}^{\prime}({\mathbb{R}})$
@@ 4.3. Boundary cohomology and its Hodge structure
By proper base change, it does not depend on the choice of the compactification $\widetilde{\mathrm{S}_{\Gamma}}$. There is a natural long exact sequence
@@
By choosing as compactification the Baily-Borel compactification $S_{\Gamma}^{*}$, mixed Hodge theory endows the spaces $\partial H^{\bullet}(\mathrm{S}_{\Gamma},V_{\lambda})$ with a canonical mixed Hodge structure, and the above long exact sequence becomes one of mixed Hodge structures.
@@
The assertion of point (1) follows from the vanishing discussed in paragraph 2.10 and from the long exact sequence (28).
To compute boundary cohomology in degree $n\in\{0,\dots,3\}$, let us first observe that since $\partial\mathrm{S}^{*}_{\Gamma}$ is 0-dimensional,
@@
Now consider the subgroup $\mathrm{M}$ of $\mathrm{T}$, and its associated Shimura datum $h_{M}$, appearing in (27). The general formalism of [27, Ch. 4] has been spelled out in this special case in [1, Lemma 3.6, Lemma 3.8]. The outcome is that the map $h_{M}$ is defined on ${\mathbb{C}}$-points by
@@
Through $h_{M}$, any representation of $\mathrm{M}_{{\mathbb{R}}}$ acquires a semisimple Hodge structure.
By [11, Thm. 2.9], there is an isomorphism of mixed Hodge structures
@@
Notice that in particular, the Hodge structure on $i_{c}^{*}R^{n}j_{*}V_{\lambda}$ is independent of the cusp $c$, so that if $m$ is the number of cusps, then as Hodge structures
@@
a computation shows that through $\chi\circ h_{M}$, the element $(z_{1},z_{2})\in{\mathbb{S}}({\mathbb{C}})$ acts on $V^{\mathrm{T}_{F}}_{\chi}$ via multiplication by
@@
So, the opposites of the exponents of $z_{1}$ and $z_{2}$ appearing above give us the type of the one-dimensional $F$-Hodge structure carried by $V^{\mathrm{T}_{F}}_{\chi}$.
@@
    The assertions in Proposition 4.1 concerning the weights of the Hodge structures of interest can also be found, with our same parametrizations, in [35, proof of Theorem 3.8]. In our way of writing the weights, we have imitated loc. cit. in emphasizing the deviation from purity (i.e., the difference, in a fixed degree $i$, between the actual weight and the pure weight $r+i$). As op. cit. shows, controlling this difference is important for being able to construct motives for cuspidal automorphic forms on $\mathrm{G}$.
@@
    The topological description of the boundary cohomology spaces $\partial H^{\bullet}(\mathrm{S}_{\Gamma},V_{\lambda})$, with no computation of their Hodge structure (since [11] was not available at that time) was already determined in [17, Eq. (2.1.2)]. In the notation of loc. cit., the elements $w_{\alpha}$, $w_{\bar{\alpha}}$, $w_{\bar{\alpha}}w_{\alpha}$, $w_{\alpha}w_{\bar{\alpha}}$ and $\theta$ correspond respectively to our (2 3)23(2\ 3)( 2 3 ), (1 2)12(1\ 2)( 1 2 ), (1 2 3)123(1\ 2\ 3)( 1 2 3 ), (1 3 2)132(1\ 3\ 2)( 1 3 2 ) and (1 3)13(1\ 3)( 1 3 ).
@@
    is an homotopy equivalence, so that the sheaf $V_{\lambda}$ canonically extends to a sheaf on $\overline{\mathrm{S}}_{\Gamma}$, denoted by the same symbol. Hence, there is no need to compute hypercohomology: the spaces $\partial H^{n}(\mathrm{S}_{\Gamma},V_{\lambda})$ are canonically isomorphic to $H^{n}(\partial\overline{\mathrm{S}}_{\Gamma},V_{\lambda})$. As a consequence, for computing boundary cohomology associated with higher-rank groups underlying a Shimura datum, the Borel-Serre compactification has a strong computational advantage over the Baily-Borel one. In our situation, the boundary structure being as simple as possible, determining boundary cohomology as vector spaces with either choice is equally easy. The Baily-Borel compactification gives us a way to describe their canonical mixed Hodge structure. In the following section, we will recall how the Borel-Serre one provides a description of boundary cohomology in terms of automorphic forms.
@@
    weight: $\mbox{\emph{weight}:}\ 0$
@@
    weight: $\mbox{\emph{weight}:}\ 1$

### v4

@@ 3.4. Bloch-Beilinson conjectures for Hecke characters
Consider the motive $M_{\phi}=(h^{1}(A_{\phi})^{\otimes w},e_{\phi})$ attached to $\phi$ as in the previous subsection. Recall that since the integer $\frac{w+1}{2}$ is critical for $M_{\phi}$, Deligne has used the comparison isomorphism between the Betti and de Rham realizations of the weight 1-1 motive $M_{\phi}(\frac{w+1}{2})$ to attach to it a canonical non-zero period $c^{+}(M_{\phi})\in{\mathbb{R}}$. He conjectured that if the value $L(M_{\phi},\frac{w+1}{2})$ does not vanish, then it should be equal to an algebraic multiple of $c^{+}(M_{\phi})$. He offered no statement when $L(\phi,s)$ has a zero at the central point. To express what is believed to happen in the latter case, consider, for a smooth, projective algebraic variety $X$ and for $n\in\{0,\dots,\dim X\}$, the subgroup $\mathop{\rm CH}\nolimits^{n}(X)_{0}$ of the Chow group $\mathop{\rm CH}\nolimits^{n}(X)$ formed by cycles homologically equivalent to 0. Bloch and Beilinson conjectured the existence of a canonical, non-degenerate height pairing
$h^{n}_{X}:\mathop{\rm CH}\nolimits^{n}(X)_{0}\otimes{\mathbb{Q}}\times\mathop{\rm CH}\nolimits^{\dim X+1-n}(X)_{0}\otimes{\mathbb{Q}}\rightarrow{\mathbb{R}}\,.$
@@
To do so, we have to extend the conjecture to the Chow groups of our motive $M_{\phi}$. Recall that for $X,n$ as above, there is a natural action of $\mathop{\rm CH}\nolimits^{\dim X}(X\times X)$ on $\mathop{\rm CH}\nolimits^{n}(X)$, preserving $\mathop{\rm CH}\nolimits^{n}(X)_{0}$. Denote by $\Gamma_{*}$ the action of an element $\Gamma\in\mathop{\rm CH}\nolimits^{\dim X}(X\times X)$. For such a $\Gamma$, the pairing $h^{n}_{X}$ restricts to a pairing, also conjecturally non-degenerate
$h^{n}_{X,\Gamma}:\Gamma_{*}\mathop{\rm CH}\nolimits^{n}(X)_{0}\otimes{\mathbb{Q}}\times\Gamma_{*}\mathop{\rm CH}\nolimits^{\dim X+1-n}(X)_{0}\otimes{\mathbb{Q}}\rightarrow{\mathbb{R}}\,.$
@@
    $\mathop{\rm\tiny{ord}}\nolimits_{s=\frac{w+1}{2}}L(\phi,s)=\dim_{{\mathbb{Q}}}(e_{\phi}e_{1,w})_{*}\mathop{\rm CH}\nolimits^{\frac{w+1}{2}}(A^{w}_{\phi})_{0}\otimes{\mathbb{Q}}$ ,
@@
    $L^{*}\left(\phi,\frac{w+1}{2}\right)\sim_{\overline{{\mathbb{Q}}}}c^{+}(M_{\phi})\det\left(h^{\frac{w+1}{2}}_{A_{\phi}^{w},e_{\phi}e_{1,w}}\right)\,.$
@@
Now, fix an element $Z\in\mathop{\rm CH}\nolimits^{\frac{w+1}{2}}(A^{w}_{\phi})_{0}$. By a completely analogous construction to the one explained in the Introduction before stating Conjecture 1 (see e.g. [32, Par. (4.2)]), such a cycle gives rise to an extension of the trivial Hodge structure by $H^{w}(A^{w}_{\phi})(\frac{w+1}{2})$. One gets an Abel-Jacobi map
$\mathop{\rm CH}\nolimits^{\frac{w+1}{2}}(A^{w}_{\phi})_{0}\rightarrow\mathop{\rm Ext}\nolimits^{1}_{\mathop{\rm MHS}\nolimits_{\overline{{\mathbb{Q}}}}}(\mathds{1},H^{w}(A^{w}_{\phi})(\frac{w+1}{2}))$
Remark 3.4.
  The name Abel-Jacobi comes from the fact that when considering mixed Hodge structures with integral coefficients, the target of the above map, by Carlson’s computation of extension spaces of mixed Hodge structures, can be identified with a complex torus (see e.g. [33, Thm. 3.31]). If we replace $A^{w}_{\phi}$ by an elliptic curve $E$ over $\mathbb{Q}$}, the space $H^{w}(A^{w}_{\phi})(\frac{w+1}{2})$ by $H^{1}(E)(1)$, and the cycle $Z$ by the divisor POP-O with $P\in E({\mathbb{Q}})$, then we find ourselves in the situation discussed in the Introduction before Conjecture 1. In this case, the complex torus in question is identified with $E({\mathbb{C}})$ itself, in such a way that the map $P\mapsto E_{P}$ can be expressed as $P\mapsto P-O$. More generally, if one replaces the elliptic curve by a higher genus smooth projective curve $C$, and POP-O by a degree 0 divisor on $C$, the complex torus is given by the complex points $J(C)({\mathbb{C}})$ of the Jacobian $J(C)$ of $C$, and the map sending a divisor to the corresponding extension is induced by the classical Abel-Jacobi map
$\mathop{\rm Div}\nolimits^{0}(C)\rightarrow J(C)({\mathbb{C}})\,.$
Passing to the subspaces cut out by the idempotent $e_{\phi}e_{1,w}$, we get
$(e_{\phi}e_{1,w})_{*}\mathop{\rm CH}\nolimits^{\frac{w+1}{2}}(A^{w}_{\phi})_{0}\rightarrow\mathop{\rm Ext}\nolimits^{1}_{\mathop{\rm MHS}\nolimits_{\overline{{\mathbb{Q}}}}}(\mathds{1},e_{\phi}e_{1,w}H^{w}(A^{w}_{\phi})(\frac{w+1}{2}))\,.$
Since $A_{\phi}^{w}$ is defined over a number field, the Abel-Jacobi map is conjectured to be injective ([8, Sect. 5.6]). Hence, the same should hold for the map above, which, using the description (29) of $H_{\phi}$, takes the form
(30) $(e_{\phi}e_{1,w})_{*}\mathop{\rm CH}\nolimits^{\frac{w+1}{2}}(A^{w}_{\phi})_{0}\rightarrow\mathop{\rm Ext}\nolimits^{1}_{\mathop{\rm MHS}\nolimits_{\overline{{\mathbb{Q}}}}}(\mathds{1},H_{\phi}(\frac{w+1}{2}))\,.$
Now recall the character $\phi^{\perp}$ defined in (28), and suppose
$\phi^{\perp}=\phi|\cdot|^{w}_{{\mathbb{I}}_{F}}\,.$
Then, by Remark 3.2, there is a well-defined sign $\epsilon(\phi)$ of the functional equation of $\phi$, and if $\epsilon(\phi)=-1$ then $L(\phi,s)$ vanishes at the central point. Hence, the conjectural injectivity of the map (30) and point (1) of Conjecture 3.3 lead to the following (considerably weaker and less precise) conjecture. In its statement, we use the terminology of geometric origin introduced at the end of Subsection 3.2.
@@
Let $\phi$ be an algebraic Hecke character of $F$ of odd weight w, such that $\phi^{\perp}=\phi|\cdot|^{w}_{{\mathbb{I}}_{F}}$. If $\epsilon(\phi)=-1$, then there exists a non-trivial extension
@@
3.5. Motives of Picard surfaces
The main aim of this paper is to construct, for a Hecke character $\phi$ verifying the required hypotheses, candidates for the extensions predicted by Conjecture 3.5. The construction will use the geometry of the Picard surfaces $\mathrm{S}_{K}$ introduced in Subsection 2.7.
More precisely, consider the universal abelian variety $A_{K}$ over $\mathrm{S}_{K}$ (Remark 2.4). For any integer $N$, we can consider its $N$-fold fibered power $A_{K}^{N}$ over $S_{K}$, and the motive $M(A^{N})$ of its total space. It is an object of $\mathop{\rm DM}\nolimits_{c}(F)$.
Fix a highest weight $\lambda$ of $\mathrm{G}_{F}$ and adopt the notations of Subsection 2.9. By the main result of [1] (see [42, before Rmk. 3.7] for a discussion in the case of Picard surfaces), there exist an integer $N$, and an idempotent endomorphism $e_{\lambda}$ of $M(A^{N})$, such that the Hodge realization of the motive
(31) $e_{\lambda}A_{K}^{N}$
coincides with $H^{\bullet}(\mathrm{S}_{K},V_{\lambda})$. This motive will be the source of the extensions of geometric origin that we want to exhibit.
@@ 4.1. Boundary of the Baily-Borel compactification
For facts recalled in this paragraph, we refer to [35, Sects. 3.6-3.7]. Any $\mathbb{Q}$}-conjugacy class of parabolics in $\mathrm{G}$ provides a set of strata of the boundary of the Baily-Borel compactification of $\mathrm{S}_{\Gamma}$. More precisely, given a representative $\mathrm{P}$ of such a conjugacy class, there is a canonical reductive subgroup $\mathrm{M}$ of the Levi component of $\mathrm{P}$, and a canonical conjugacy class of morphisms
@@ 4.2. Boundary of the Borel-Serre compactification
For facts recalled in this paragraph, we refer to [23, pp. 567-570]. The Borel-Serre compactification $\overline{\mathrm{S}}_{\Gamma}$ of $\mathrm{S}_{\Gamma}$ admits a canonical projection $\pi$ to the Baily-Borel one. Let us describe, in our situation, the fiber of $\pi$ over each cusp $p$ of $\mathrm{S}^{*}_{\Gamma}$. To do so, let us write
@@
Denote by $A_{\mathrm{G}}$ the identity component of the real points of the maximal $\mathbb{Q}$}-split subtorus $\mathrm{T}^{s}$ of $\mathrm{G}^{\mathop{\rm\tiny{der}}\nolimits}$. From (11), we see that there is an isomorphism of Lie groups $\mathrm{T}^{\prime}({\mathbb{R}})\simeq{\mathbb{C}}^{\times}$ which by (12) restricts to $A_{\mathrm{G}}\simeq{\mathbb{R}}^{\times}_{>0}.$
[…]
@@ 4.3. Boundary cohomology and its Hodge structure
where $Rj_{*}$ stands for the total derived direct image functor with respect to $j$.
By proper base change, boundary cohomology does not depend on the choice of the compactification $\widetilde{\mathrm{S}_{\Gamma}}$. There is a natural long exact sequence
@@
and a canonical mixed Hodge structure on the spaces $\partial H^{\bullet}(\mathrm{S}_{\Gamma},V_{\lambda})$ making the above a long exact sequence of mixed Hodge structures. This can be seen, for example, by applying the formalism of mixed Hodge modules mentioned in Subsection 2.9.
@@
The assertion of point (1) follows from the vanishing discussed in Subsection 2.10 and from the long exact sequence (33).
To compute boundary cohomology in degree $n\in\{0,\dots,3\}$, we choose to use the Baily-Borel compactification $\mathrm{S}^{*}_{\Gamma}$. Then, since $\partial\mathrm{S}^{*}_{\Gamma}$ is 0-dimensional,
@@
Now consider the subgroup $\mathrm{M}$ of $\mathrm{T}$, and its associated Shimura datum $h_{\mathrm{M}}$, appearing in (32). The general formalism of [34, Ch. 4] has been spelled out in this special case in [2, Lem. 3.6, Lem. 3.8]. The outcome is that the map $h_{M}$ is defined on $\mathbb{C}$}-points by
@@
Through $h_{\mathrm{M}}$, any representation of $\mathrm{M}_{{\mathbb{R}}}$ acquires a semisimple Hodge structure.
By [15, Thm. 2.9], there is an isomorphism of mixed Hodge structures
@@
In particular, notice that the Hodge structure on $i_{c}^{*}R^{n}j_{*}V_{\lambda}$ is independent of the cusp $c$, so that if $m$ is the number of cusps, then as Hodge structures
@@
a computation shows that through $\chi\circ h_{\mathrm{M}}$, the element $(z_{1},z_{2})\in{\mathbb{S}}({\mathbb{C}})$ acts on $V^{\mathrm{T}_{F}}_{\chi}$ via multiplication by
@@
So, the negatives of the exponents of $z_{1}$ and $z_{2}$ appearing above give us the type of the one-dimensional $F$-Hodge structure carried by $V^{\mathrm{T}_{F}}_{\chi}$.
@@
    The assertions in Proposition 4.1 concerning the weights of the Hodge structures of interest can also be found, with our same parametrizations, in [42, proof of Thm. 3.8]. In our way of writing the weights, we have imitated loc. cit. in emphasizing the deviation from purity (i.e., the difference, in a fixed degree $i$, between the actual weight and the pure weight $r+i$). As op. cit. shows, controlling this difference is important for being able to construct motives for cuspidal automorphic forms on $\mathrm{G}$.
@@
    The topological description of the boundary cohomology spaces $\partial H^{\bullet}(\mathrm{S}_{\Gamma},V_{\lambda})$, with no computation of their Hodge structure (since [15] was not available at that time) was already determined in [23, Eq. (2.1.2)]. In the notation of loc. cit., the Weyl elements $w_{\alpha}$, $w_{\bar{\alpha}}$, $w_{\bar{\alpha}}w_{\alpha}$, $w_{\alpha}w_{\bar{\alpha}}$ and $\theta$ correspond respectively to our elements (2 3)(2\ 3), (1 2)(1\ 2), (1 2 3)(1\ 2\ 3), (1 3 2)(1\ 3\ 2) and (1 3)(1\ 3).
@@
    is an homotopy equivalence, so that the sheaf $V_{\lambda}$ canonically extends to a sheaf on $\overline{\mathrm{S}}_{\Gamma}$, denoted by the same symbol. Hence, there is no need to compute hypercohomology: the spaces $\partial H^{n}(\mathrm{S}_{\Gamma},V_{\lambda})$ are canonically isomorphic to $H^{n}(\partial\overline{\mathrm{S}}_{\Gamma},V_{\lambda})$. As a consequence, for computing boundary cohomology associated with higher-rank groups underlying a Shimura datum, the Borel-Serre compactification has a strong computational advantage over the Baily-Borel one. In our situation, the boundary structure being as simple as possible, determining boundary cohomology as vector spaces with either choice is equally easy. The Baily-Borel compactification gives us a way to describe their canonical mixed Hodge structure. In the following section, we will recall how the Borel-Serre one provides a description of the $\mathrm{G}({\mathbb{A}}_{f})$-module structure of boundary cohomology.
@@
    $\mbox{\emph{weight}:}\ 0$
@@
    $\mbox{\emph{weight}:}\ 1$